A hotel claims that 85% of its customers are very satisfied with its service. Complete parts a through d below based on a ranThe answers are in red. I just want to know how to correctly solve this...
Answer #1

- Ans: P = 0.85, n = 5 w xn Bm (n=5, P=0.85) P (x = 4) - 5 (0.85) (0.15) = 10.3915] 6 p (x34) = P(x-5) = 5 (0.85 (0.150 = (0